What is the magnitude of the electric force between two electrons

separated by a distance of 0.10 nm? (1 nm = 1 x 10 -9 m)


1
Expert's answer
2022-02-25T11:14:52-0500

"F=\frac{kq^2}{r^2}=2.3\cdot 10^{-8}~N."